OpenStudy (anonymous):

FODQ, solve for general solution (x-y)y'=x+y y' =(x+y)/(x-y) y'/x = (1+y/x)/(1-y/x) let v = y/x v/x + v' = (1+v)/(1-v) it appears I can't cleanly seperate these variables, but this is clearly homogeneous..

OpenStudy (anonymous):

y = 2tan(y/x)^-1 - ln(y^2/x^2 + 1) = 2ln(x) + c

OpenStudy (anonymous):

good way to test as to is it homogeneous....given f' = (x,y), then f(x, xt) must equal f(1,t)

OpenStudy (anonymous):

never heard of such a test, I'm interested. It seemed most likely in this form compared to direct or bernoulli

OpenStudy (anonymous):

all homogeneous first order d.eq's are separable after v=y/x substi.

OpenStudy (anonymous):

so I was making lemonade out of charcoal? then this must be direct? I've attemped direct and could't seperate them either

OpenStudy (anonymous):

ayan -try my cute definition for homogeneous 1st orders

OpenStudy (anonymous):

v =y/x vx=y y'= v+v'x

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

i don't follow however, i've used those subsitutions, and ended up with an unseperable equation

OpenStudy (anonymous):

all homogenous 1st orders can be separated

OpenStudy (anonymous):

am I utilizing the wrong subsitution?

OpenStudy (anonymous):

I can't see bernoulli will help me in this situation.

OpenStudy (anonymous):

bernoulli pf of the fork y'+ f(x)y = g(x) * y^n where n cannot = 0 or 1 pf should read is...fork reads form

OpenStudy (anonymous):

i teach diff eq. using boyce and diprima's brilliant book

OpenStudy (anonymous):

thank you, I haven't solved this yet, but i'll try hacking at it my putting it in bernoulli form. Given the daunting final result, I'm not sure if I'll calculate it correctly.

OpenStudy (anonymous):

what is the problem?

OpenStudy (anonymous):

hacking into it again now.

OpenStudy (anonymous):

your sub is wrong...the left side should be v + V'x

OpenStudy (anonymous):

y' = v + v'x since i have y'/x this equates to v/x + v' does it not?

OpenStudy (anonymous):

you have y' = (x+y)/(x-y) = v+v'x.....times the left by i/x over 1/x, and iys homogeneous and ready for its algebra helper

OpenStudy (anonymous):

(1+v)/ (1-v) - v =v'x...that separates

OpenStudy (anonymous):

where v' is dv/dx

OpenStudy (anonymous):

after cleaning up the rational expression on the left

OpenStudy (anonymous):

I'm probably viewing my problem thought a kolidascope, but I don't see how you disposed of y'/x = (1+y/x)/(1-y/x) prior to subsitution

OpenStudy (anonymous):

where both sides needed to be divided by x, sorry if I'm sounding imbecile

OpenStudy (anonymous):

from there, my left side after subsitution becomes (v+v'x)/x

OpenStudy (anonymous):

ur prob was y' = (x+y)/(x-y) y' = v+v'x = (1+v)/(1-v)....subtract v from both sides and get a common denominator so, (1+v)/1-v) - v/1= 1+v -(1-v)*v/1 all over 1-v

OpenStudy (anonymous):

(1+v-v+v^2) = 1+v^2 / 1-v = v'x

OpenStudy (anonymous):

(1-v)dv / (1+v^2) =dx/x

OpenStudy (anonymous):

i'm still rolling around the first statement in my head, I under stand how you proceed from y'=v+v'x = (1+v)/(1-v) it's just that to get that v sub in, I still believe that y'/x had to have occured. Why do I believe that I wonder

OpenStudy (anonymous):

v=y/x.......to get y'. you need y = vx

OpenStudy (anonymous):

then product rule with vx

OpenStudy (anonymous):

yes, y' = x'v + v'x and x' = 1 so v'=v + v'x I'm shaking the stupid from my head but this nugget is stuborn

OpenStudy (anonymous):

but i have only taught this stuff for 20 yrs :)

OpenStudy (anonymous):

no no y' = v +v'x

OpenStudy (anonymous):

I must be a vast disapointmet to encounter then by your reckoning..

OpenStudy (anonymous):

yes y', sorry, typo on that occurance

OpenStudy (anonymous):

not at all...we are all tiny players

OpenStudy (anonymous):

do you see how it separates now?

OpenStudy (anonymous):

if i could do it on a board for you, it would take 2 minutes

OpenStudy (anonymous):

yes but I still have that nugget, I think I'm thinking something wrong. In all honesty, the integral is supposed to be harder than seperating, I have no idea why this bump is becoming a boulder

OpenStudy (anonymous):

all homo first orders can be separated

OpenStudy (anonymous):

so I see it, I'm thinking, that i'm just doing a mathematical error similar to losing a negative of sorts involving 1/x

OpenStudy (anonymous):

perhaps I should state my problem clearly

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[y' = (x+y)/(x-y)\] mult 1/x on LHS and RHS \[y'/x = (x/x + y/x)/(x/x - y/x) = (1+ y/x)(1-y/x)\] let v = y/x so \[y'/x = (1+v)/(1-v)\] = \[(v+v'x)/x = (1+v)/(1-v)\]

OpenStudy (anonymous):

i suppose you can see my rouge 1/x , probably a product of my stupidity

OpenStudy (anonymous):

after this however, I can cleanly integrate it, I suppose I'm just not doing my basic fractions right

OpenStudy (anonymous):

you only times the top nad bottom by 1/x not the other side....y' reamins as is v + v'x

OpenStudy (anonymous):

oh, yes.. and it doesn't change the equation as it's anagolous to rewriting the equation

OpenStudy (anonymous):

I feel very stupid now

OpenStudy (anonymous):

v+v'x = (1+v)/(1-v)

OpenStudy (anonymous):

yes, I see how to proceed, how did I let that slip. Thank you for letting me waste your time, and I appologize for that. Apparently I'm losing grip with my sanity.